# On Large-scale Evaluation of Embedding Models for Knowledge Graph Completion

### v1

@@ 3.2 Global Metrics
4 Problems of the Generic Evaluation Protocol
This section examines the limitations of link prediction as a widely used evaluation protocol for assessing KGE models. We present the results from alternative evaluation protocols, including entity-pair ranking, property prediction, and triple classification. Our results indicate that the models perform poorly on these alternative evaluations and there is some mismatch between link prediction accuracy and performance on other tasks. For instance, while RotatE achieves the highest link prediction scores, DistMult and ComplEx excel in triple classification. These findings highlight the need for better training and evaluation strategies for KGE models.
4.1 Link Prediction
@@
4.2 Entity-Pair Ranking
Link prediction evaluation setup can be misleading as we cannot verify if a model ranks all types of false or nonsensical triples lower than correct triples. For example, consider a test triple (James Ivory, director/film, A Room With A View). Evaluation by link prediction only seeks responses to two sensible questions: “who directed the film A Room With A View?”, i.e., (?, director/film, A Room With A View), and “which film(s) are directed by James Ivory?”, i.e., (James Ivory, director/film, ?). It does not verify if a model would rank false or nonsensical triples such as (A Room With A View, director/film, James Ivory) lower than the correct one. In this triple, the subject and object entities are swapped, resulting in an incorrect direction for the relation. Wang et al. [40] discussed this problem and proposed to use an alternative protocol called entity-pair ranking. Following this evaluation protocol, we would create all possible combinations of entities for a relation and then rank them based on their scores. Specifically, Wang et al. [40] proposed using Weighted Mean Average Precision (WMAP@K) and Weighted Precision (WP@K). To treat each relation equally, we employ macro-averaging, which corresponds to reporting the unweighted version of these metrics.
Specifically, the unweighted version of $WMAP@K$ is defined as follows:
$MAP@K=\sum_{r\in\mathcal{R}}AP_{r}@K$ (1)
where $R$ is the set of all relations in the test set and $AP_{r}@K$ is defined as follows:
$AP_{r}@K=\frac{1}{NP_{r}}\sum_{k\in{K}}P_{r}@k\times relv_{r}[k]$ (2)
where $NP_{r}$ is the lesser of $K$ and $|T_{r}|$ (the number of test triples for relation $r$), $P_{r}@k$ is the precision among top $k$ predictions for relation $r$, and $relv_{r}[k]$ is a Boolean indicator of whether the $k$th prediction is correct (1) or not (0).
Furthermore, we define the unweighted version of $WP@K$ as follows:
$P@K=\sum_{r\in\mathcal{R}}P_{r}@K$ (3)
where $R$ is the set of all relations in the test set and $P_{r}@K$ is the precision among top $K$ predictions for relation $r$.
@@ 5 Conclusion
5.0.1 Acknowledgements
This material is based upon work supported by the National Science Foundation under Grants TIP-2333834. We also extend our gratitude to the Texas Advanced Computing Center (TACC) for providing compute resources used in this work’s experimentation.

### v2

@@ 5.1.5 Macro-Averaging by Domain
5.2 Mitigating Weaknesses in Evaluation Protocols
In this section, we reproduce link prediction results and present the results from alternative evaluation protocols, including entity-pair ranking, property prediction, and triple classification. Our results indicate that the models perform poorly on these alternative evaluations, and there is some mismatch between link prediction accuracy and performance on other tasks. For instance, while RotatE achieves the highest link prediction scores, DistMult and ComplEx excel in triple classification. These findings highlight the need for better training and evaluation strategies for KGE models.
5.2.1 Link Prediction
@@
5.2.2 Entity-Pair Ranking
Wang et al. [40] discussed the first limitation with link prediction, mentioned in Section 4.2 and proposed to use an alternative protocol called entity-pair ranking. Following this evaluation protocol, we would create all possible combinations of entities for a relation and then rank them based on their scores. Specifically, Wang et al. [40] proposed using Weighted Mean Average Precision (WMAP@K) and Weighted Precision (WP@K). To treat each relation equally, we employ macro-averaging, which corresponds to reporting the unweighted version of these metrics.
Specifically, the unweighted version of $WMAP@K$ is defined as $MAP@K=\sum_{r\in\mathcal{R}}AP_{r}@K$, where $R$ is the set of all relations in the test set and $AP_{r}@K$ is defined as $AP_{r}@K=\frac{1}{NP_{r}}\sum_{k\in{K}}P_{r}@k\times relv_{r}[k]$ in which $NP_{r}$ is the lesser of $K$ and $|T_{r}|$ (the number of test triples for relation $r$), $P_{r}@k$ is the precision among top $k$ predictions for relation $r$, and $relv_{r}[k]$ is a Boolean indicator of whether the $k$th prediction is correct (1) or not (0). Furthermore, we define the unweighted version of $WP@K$ as $P@K=\sum_{r\in\mathcal{R}}P_{r}@K$ where $R$ is the set of all relations in the test set and $P_{r}@K$ is the precision among top $K$ predictions for relation $r$.